Write the equation of the line that passes through point (–1, –3) with a slope of –5.
Answer by checkley75(3666)

Y=mX+b WHERE m=SLOPE & b=Y INTERCEPT.
NOW REPLACE X & Y WITH THE POINT(-1,-3) ANS USING A SLOPE OF -5 WE CAN FIND THE Y INTERCEPT(b).
-3=-5*-1+b
-3=5+b
b=-3-5
b=-8 THE Y INTERCEPT.
SO THE LINE EQUATION IS: Y=-5X-8